OpenStudy (anonymous):

Associatative law of addition to write an equivalent expression for (x+3a)+y

OpenStudy (anonymous):

The associative property states that you may perform different additions first. So an equivalent expression would be x + (3a+y)

Directrix (directrix):

Associate Property for Addition is a grouping property. It states that for numbers a, b, and c, a + (b + c) = (a +b) + c. For, x + (3a+y), the Associative Property would say x + (3a+y) = (x + 3a) + y. Notice that the x and 3a are now grouped and the y is "solo" whereas before the 3a and y were grouped and the x was "solo."